Diamond Crystal

Given an integer value 

N, write a program to print a diamond pattern of 2*N rows as shown below.

Input

The first line of input is an integer 

N.

Explanation

In the given example, the number of rows in the diamond is 

2*5 = 10.

So, the output should be




    /\
   /  \
  /    \
 /      \
/        \
\        /
 \      /
  \    /
   \  /
    \/
1
Expert's answer
2021-09-12T11:48:57-0400
N = int(input())
for i in range(N):
    print(' ' * (N - i - 1), end='')
    print('/', end='')
    print(' ' * (2 * i), end='')
    print('\\')
for i in range(N - 1, -1, -1):
    print(' ' * (N - i - 1), end='')
    print('\\', end='')
    print(' ' * (2 * i), end='')
    print('/')